# S-box Security Analysis of NIST Lightweight Cryptography Candidates: A Critical Empirical Study

### v1

@@ 6 Side Channel Analysis
$8\times 8$ Sbox
Ideal Value
1
$\mathsf{OP}$
$1\leq\mathsf{OP}\leq 2^{n}$ $1\leq\mathsf{OP}\leq 16$ $1\leq\mathsf{OP}\leq 32$ $1\leq\mathsf{OP}\leq 256$ Close to LB
2
$\mathsf{FP}$
$0\leq\mathsf{FP}\leq 2^{n}$ $0\leq\mathsf{FP}\leq 16$ $0\leq\mathsf{FP}\leq 32$ $0\leq\mathsf{FP}\leq 256$ Close to LB
3
$\mathsf{OFP}$
$0\leq\mathsf{OFP}\leq 2^{n}$ $0\leq\mathsf{OFP}\leq 16$ $0\leq\mathsf{OFP}\leq 32$ $0\leq\mathsf{OFP}\leq 256$ Close to LB
4
$\mathsf{BIC}$
$0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ 00
5
$\mathsf{SAC}$
$0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0.5$
6
$\barwedge_{S}$
$\sqrt{{\frac{2^{2n}}{2^{n}-1}}}\leq\barwedge_{S}\leq 2^{n}$
$0<\barwedge_{S}\leq 16$
$0<\barwedge_{S}\leq 32$
$0<\barwedge_{S}\leq 256$
Close to LB
7
$\doublebarwedge_{S}$
$2^{2n}\leq\doublebarwedge_{S}\leq 2^{(3n+m)}$ $256\leq\doublebarwedge_{S}\leq 65536$ $1024\leq\doublebarwedge_{S}\leq 1048576$ $65536\leq\doublebarwedge_{S}\leq 4294967296$ Close to LB
8
$\mathsf{LAT}$
$0\leq\mathsf{LAT}\leq 2^{n-1}$ $0\leq\mathsf{LAT}\leq 8$ $0\leq\mathsf{LAT}\leq 16$ $0\leq\mathsf{LAT}\leq 128$ Close to LB
9
$\xi$
$0\leq\xi\leq 2^{-1}$ $0\leq\xi\leq 2^{-1}$ $0\leq\xi\leq 2^{-1}$ $0\leq\xi\leq 2^{-1}$ Close to LB
10
$\mathsf{NL}$
For $n:even$
$0\leq\mathsf{NL}<2^{n-1}-2^{\frac{n}{2}-1}$
For $n:odd$
$0\leq\mathsf{NL}<2^{n-1}-2^{\frac{n-1}{2}}$
$60\leq\mathsf{NL}<$ 6 $120\leq\mathsf{NL}<$ 12 $1200\leq\mathsf{NL}<$ 120 Close to UB
11
$\mathsf{BN_{L}}$
$12\leq\mathsf{BN_{L}}\leq n-$ 1 $2\leq\mathsf{BN_{L}}\leq 3$ $2\leq\mathsf{BN_{L}}\leq 4$ $2\leq\mathsf{BN_{L}}\leq 7$ Close to UB
12
$\mathsf{LS}$
$0\leq\mathsf{LS}\leq 2^{n+m}$ $0\leq\mathsf{LS}\leq 256$ $0\leq\mathsf{LS}\leq 1024$ $0\leq\mathsf{LS}\leq 65536$ Close to LB
13
$\mathsf{CI}$
$0\leq\mathsf{CI}\leq n$ $0\leq\mathsf{CI}\leq 4$ $0\leq\mathsf{CI}\leq 5$ $0\leq\mathsf{CI}\leq 8$ Close to LB
14
$\mathsf{DU}$
$2\leq\mathsf{DU}\leq 2^{n}$ $2\leq\mathsf{DU}\leq 16$ $2\leq\mathsf{DU}\leq 32$ $2\leq\mathsf{DU}\leq 256$ Close to LB
15
$\mathsf{BN_{D}}$
$2\leq\mathsf{BN_{D}}\leq\lceil{\frac{2n}{3}}\rceil$ $2\leq\mathsf{BN_{L}}\leq 3$ $2\leq\mathsf{BN_{D}}\leq 4$ $2\leq\mathsf{BN_{D}}\leq 6$ Close to UB
16
$\mathsf{PC}$
$0\leq\mathsf{PC}\leq n$  $0\leq\mathsf{PC}\leq 4$ $0\leq\mathsf{PC}\leq 5$ $0\leq\mathsf{PC}\leq 8$ Close to UB
17
$\mathsf{UDB}$
$\mathsf{UDB}=0$ $\mathsf{UDB}=0$ $\mathsf{UDB}=0$ $\mathsf{UDB}=0$ 00
18
$\beta\mathsf{U}$
$2\leq\beta\mathsf{U}\leq 2^{n}$ $2\leq\beta\mathsf{U}\leq 16$ $2\leq\beta\mathsf{U}\leq 32$ $2\leq\beta\mathsf{U}\leq 256$ Close to LB
19
$\mathsf{DLU}$
$2^{\frac{n}{2}-1}<\mathsf{DLU}\leq 2^{n-1}$ $2<\mathsf{DLU}\leq 8$ $4<\mathsf{DLU}\leq 16$ $8<\mathsf{DLU}\leq 128$ Close to LB
20
$\mathsf{AD}$
$11\leq\mathsf{AD}\leq n-$ 1 $1\leq\mathsf{AD}\leq 3$ $1\leq\mathsf{AD}\leq 4$ $1\leq\mathsf{AD}\leq 7$ Close to UB
21
$\mathsf{DPA_{SNR}}$
@@
$1\leq\mathsf{DPA_{SNR}}\leq 2^{\frac{n}{2}}$
$0\leq\mathsf{DPA_{SNR}}\leq 4$ $60\leq\mathsf{DPA_{SNR}}<$ 6 $0\leq\mathsf{DPA_{SNR}}\leq 16$ Close to LB
22
$\mathsf{TO}$
$0\leq\mathsf{TO}\leq m$ $0\leq\mathsf{TO}\leq 4$ $0\leq\mathsf{TO}\leq 5$ $0\leq\mathsf{TO}\leq 8$ Close to LB
3 Cryptographic Properties of Sboxes and their Classification
The strength of cryptographic primitives is best analysed with the help of well known cryptanalysis techniques. Shannon’s theory of confusion and diffusion plays the foundational role in the design and analysis of iterative ciphers. The aspect of confusion is introduced in symmetric ciphers using highly non-linear components known as sboxes. Designers and cryptanalysts treat these ciphers as a white box, analysing all the underlying mathematical components in order to understand and rule out any potential attack vectors, by imposing design practices based on standard characteristics and theoretically well-defined bounds. Owing to the importance of sboxes in the design of cryptographic algorithms, an extensive research has been done over the years to explain the most refined design and analysis criteria that could be employed in order to secure them against potential attack vectors. These properties can be classified in a number of ways depending whether upon the types of representation in boolean algebra or the statistical and implementation based attacks that could be launched through exploitation of relevant metrics or via studying the inter-dependencies that link one property to another.
This section briefly covers an introduction to around twenty-five cryptographic properties of sboxes that hold significance in the current literature. Each of the properties is related to one or more crytanalysis techniques, that could exploit the non-linear part of a block cipher, see Table 2. In addition, Table 3 presents the upper and lower bounds (UB & LB) on these properties with respect to the dimension of sbox being used [74, 75, 59, 76, 77, 78, 60, 61]. We also talk about the behaviour of these properties under affine equivalence, see Table 4. The properties along with their respective bounds and equivalence characteristics aid designers and cryptanalysts in the selection of good sboxes.
3.1 Diffusion Characteristics
Properties such as balancedness, bijectivity, order of permutation, fixed and opposite fixed points, strict avalanche and bit independence criteria, absolute and sum of square indicator are general pre-requisites that should adhere to defined bounds for any sbox in order to avoid correlations between the input and output bits and serve as a pre-requisite to ensure security against various statistical cryptanalysis attacks.
Balancedness
A vectorial boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ is balanced if it has an equal number of $1$’s and 00’s in its coordinate and component functions. Mathematically;
$P(0)=P(1)=\frac{1}{2}=50\%$ (1)
An imbalanced boolean mapping stays at a risk of achieving high linear approximation probabilities [79]. Balancedness is invariant when under affine equivalence, which means that the number of 00’s ad $1$’s is preserved in $S$ if an affine transformation is applied to it [80, 81, 82].
Permutation
Order of Permutation (OP)
Fixed Points (FP)
$S(\varkappa)=\varkappa$ (2)
Opposite Fixed Points (OFP)
$S(\varkappa)=({\varkappa})^{\prime}$ (3)
Bit Independence Criteria (BIC)
A vectorial boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ is said to satisfy BIC, if inverting an input bit $u$ has a changing effect on the output bit $v$ i.e. all the avalanche variables $(\gamma_{u},\gamma_{v},....)$ representing difference between bits of input and output of $S$ become pairwise independent [85]. Mathematically, BIC is computed using correlation co-efficient of the avalanche variables:-
[…]
@@ 3.1 Diffusion Characteristics
For the coordinate and component functions of $S$ the BIC value should stay in the range of $[0,1]$. An ideal sbox $S$ has a BIC value equal to 00. If an affine transformation is applied to $S$, BIC does not stay invariant [86].
@@
Where $\gamma\in F_{2}^{n},\gamma\neq 0,wt(\gamma)=1$ SAC is variant under affine transformation, which means that a function not satisfying SAC could be transformed to a function that satisfies SAC using affine equivalence [87].
@@
The overall impact of auto correlation of a boolean function or a vectorial boolean function is studied by finding the maximum value in ACT spectrum, known as absolute indicator (AI).
Let $f$ be a Boolean function over $F_{2}^{n}$ i.e. $f:F_{2}^{n}\rightarrow F_{2}$. The AI of $f$ is defined as [74]:-
@@
Here $D_{\gamma}f=f(x)\oplus f(x\oplus a)$.
@@
$\barwedge_{S}$ stays invariant under affine transformation. For balanced vectorial boolean functions, the value of AI is always 00. If a boolean mapping has high $\barwedge_{S}$, it becomes highly susceptible to cube attacks [89, 74].
@@
SSI is not variant under affine equivalence [90]. For achieving good diffusion characteristics in a cryptographic algorithm, sboxes should have low SSI and AI values [91, 74].
@@ 3.2 Linear Cryptanalysis
The idea of this technique evolves around studying the non-linear layer of a block cipher to draw a linear approximations profile (LAT) between plaintext, key and the ciphertext to find correlations among them. Matsui introduced non-linearity as a significant criteria in the design of block ciphers [92]. Block ciphers having a highly non-linear layer are resistant towards linear cryptanalysis. To study the number of active sboxes in the linear trails of a block cipher, J. Daemon et al. [93] introduced linear branch number. Theoretically, in order to resist linear (and its variant) cryptanalysis techniques an sbox should have a low linear approximation probability, high non-linearity, high order of linear branch number, equal to none linear structures, and low order correlation immunity.
@@
$\mathsf{LAT}(\varrho,\psi)=\{\#\varkappa\ |\ \varrho\cdot\varkappa=\psi\cdot S(\varkappa)\}$ (11)
Definition 3.1
@@
$\xi=\frac{1}{2^{n}}\max_{\varrho,\psi\in F_{2}^{n},F_{2}^{m}\backslash\{0\}}\bigg{|}\sum_{\varkappa=1}^{2^{n-1}}\mathsf{LAT}(\varrho,\psi)\bigg{|}$ (12)
@@
In block cipher design strategy, branch number is a significant property used to measure the strength of an sbox against linear cryptanalysis. An sbox with high branch number could increase the diffusion power of a cryptographic algorithm. Mathematically:-
@@
LBN is variant under affine transformation i.e. two sboxes can be affine equivalent of each other but they can have different values of LBN [98].
@@
Sboxes having LS are vulnerable to attacks more robust than exhaustive key search, such as differential and linear cryptanalysis [100]. LS is an affine variant property of any boolean mapping $S$ [101, 74].
@@
[…]
CI shows variant behaviour when an affine transformation is applied to $S$ [74]. For a good cryptographic sbox the lower is the order of CI the better. CI shares an inverse relation with NL, hence if CI holds a lower order the value of NL will be high, and vice versa.

### v3

@@ 7 Side-Channel Analysis DPA-SNR [72]; Transparency Order (TO) [73]
8×8
Ideal
Value
1 $\mathsf{OP}$ $1\leq\mathsf{OP}\leq 2^{n}$ $1\leq\mathsf{OP}\leq 16$ $1\leq\mathsf{OP}\leq 32$ $1\leq\mathsf{OP}\leq 256$ Close to UB
2 $\mathsf{FP}$ $0\leq\mathsf{FP}\leq 2^{n}$ $0\leq\mathsf{FP}\leq 16$ $0\leq\mathsf{FP}\leq 32$ $0\leq\mathsf{FP}\leq 256$ Close to LB
3 $\mathsf{OFP}$ $0\leq\mathsf{OFP}\leq 2^{n}$ $0\leq\mathsf{OFP}\leq 16$ $0\leq\mathsf{OFP}\leq 32$ $0\leq\mathsf{OFP}\leq 256$ Close to LB
4 $\mathsf{BIC}$ $0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ $0\leq\mathsf{BIC}\leq 1$ 0
5 $\mathsf{SAC}$ $0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0\leq\mathsf{SAC}\leq 1$ $0.5$
6 $\barwedge_{S}$ $\sqrt{\tfrac{2^{2n}}{2^{n}-1}}\leq\barwedge_{S}\leq 2^{n}$ $0<\barwedge_{S}\leq 16$ $0<\barwedge_{S}\leq 32$ $0<\barwedge_{S}\leq 256$ Close to LB
7 $\doublebarwedge_{S}$ $2^{2n}\leq\doublebarwedge_{S}\leq 2^{3n+m}$ $256\leq\doublebarwedge_{S}\leq 65536$ $1024\leq\doublebarwedge_{S}\leq 1048576$ $65536\leq\doublebarwedge_{S}\leq 4294967296$ Close to LB
8 $\mathsf{LAT}$ $0\leq\mathsf{LAT}\leq 2^{n}-1$ $0\leq\mathsf{LAT}\leq 15$ $0\leq\mathsf{LAT}\leq 31$ $0\leq\mathsf{LAT}\leq 255$ -
9 $\xi$ $0\leq\xi\leq\tfrac{1}{2}$ $0\leq\xi\leq\tfrac{1}{2}$ $0\leq\xi\leq\tfrac{1}{2}$ $0\leq\xi\leq\tfrac{1}{2}$ Close to LB
10 $\mathsf{NL}$
For $n$ even: $0\leq\mathsf{NL}<2^{n-1}-2^{\tfrac{n}{2}-1}$
For $n$ odd: $0\leq\mathsf{NL}<2^{n-1}-2^{\tfrac{n-1}{2}}$
$0\leq\mathsf{NL}<6$ $0\leq\mathsf{NL}<12$ $0\leq\mathsf{NL}<120$ Close to UB
11 $\mathsf{BN_{L}}$ $2\leq\mathsf{BN_{L}}\leq n-1$ $2\leq\mathsf{BN_{L}}\leq 3$ $2\leq\mathsf{BN_{L}}\leq 4$ $2\leq\mathsf{BN_{L}}\leq 7$ Close to UB
12 $\mathsf{LS}$ $0\leq\mathsf{LS}\leq 2^{n+m}$ $0\leq\mathsf{LS}\leq 256$ $0\leq\mathsf{LS}\leq 1024$ $0\leq\mathsf{LS}\leq 65536$ Close to LB
13 $\mathsf{CI}$ $0\leq\mathsf{CI}\leq n$ $0\leq\mathsf{CI}\leq 4$ $0\leq\mathsf{CI}\leq 5$ $0\leq\mathsf{CI}\leq 8$ Close to LB
14 $\mathsf{DDT}$ $0\leq\mathsf{DDT}\leq 2^{n}-1$ $0\leq\mathsf{DDT}\leq 15$ $0\leq\mathsf{DDT}\leq 31$ $0\leq\mathsf{DDT}\leq 255$ -
15 $\mathsf{DU}$ $2\leq\mathsf{DU}\leq 2^{n}$ $2\leq\mathsf{DU}\leq 16$ $2\leq\mathsf{DU}\leq 32$ $2\leq\mathsf{DU}\leq 256$ Close to LB
16 $\mathsf{BN_{D}}$ $2\leq\mathsf{BN_{D}}\leq\lceil\tfrac{2n}{3}\rceil$ $2\leq\mathsf{BN_{D}}\leq 3$ $2\leq\mathsf{BN_{D}}\leq 4$ $2\leq\mathsf{BN_{D}}\leq 6$ Close to UB
17 $\mathsf{PC}$ $0\leq\mathsf{PC}\leq n$ $0\leq\mathsf{PC}\leq 4$ $0\leq\mathsf{PC}\leq 5$ $0\leq\mathsf{PC}\leq 8$ Close to UB
18 $\mathsf{UDB}$ $\mathsf{UDB}=0$ $\mathsf{UDB}=0$ $\mathsf{UDB}=0$ $\mathsf{UDB}=0$ 0
19 $\beta\mathsf{U}$ $2\leq\beta\mathsf{U}\leq 2^{n}$ $2\leq\beta\mathsf{U}\leq 16$ $2\leq\beta\mathsf{U}\leq 32$ $2\leq\beta\mathsf{U}\leq 256$ Close to LB
20 $\mathsf{DLU}$ $2^{\tfrac{n}{2}-1}<\mathsf{DLU}\leq 2^{n-1}$ $2<\mathsf{DLU}\leq 8$ $4<\mathsf{DLU}\leq 16$ $8<\mathsf{DLU}\leq 128$ Close to LB
21 $\mathsf{AD}$ $1\leq\mathsf{AD}\leq n-1$ $1\leq\mathsf{AD}\leq 3$ $1\leq\mathsf{AD}\leq 4$ $1\leq\mathsf{AD}\leq 7$ Close to UB
22 $\mathsf{DPA_{SNR}}$
@@
$1\leq\mathsf{DPA_{SNR}}\leq 2^{\tfrac{n}{2}}$
$0\leq\mathsf{DPA_{SNR}}\leq 4$ $0\leq\mathsf{DPA_{SNR}}<6$ $0\leq\mathsf{DPA_{SNR}}\leq 16$ Close to LB
23 $\mathsf{TO}$ $0\leq\mathsf{TO}\leq m$ $0\leq\mathsf{TO}\leq 4$ $0\leq\mathsf{TO}\leq 5$ $0\leq\mathsf{TO}\leq 8$ Close to LB
@@
[…]
Table 6: Results of S-box Analysis – $8\times 8$ and $5\times 5$ bit case
  S# Properties
Romulus
$8\times 8$ Skinny-128
ASCON/ISAP
$5\times 5$
  1 Bijectivity
  2 Balancedness
  3 Permutation
  4 $\mathsf{OP}$ 140 26
  5 $\mathsf{FP}$ 1 0
  6 $\mathsf{OFP}$ 0 0
  7 $\mathsf{BIC}$ 1.000000 1.000000
  8 $\mathsf{SAC}$ 1.000000 1.000000
  9 $\barwedge_{S}$ 256 32
  10 $\doublebarwedge_{S}$ 4194304 8192
  11 $\xi$ 0.25000 0.25000
  12 $\mathsf{NL}$ 64 8
  13 $\mathsf{BN_{L}}$ 2 3
  14 $\mathsf{LS}$ 601 91
  15 $\mathsf{CI}$ 0 0
  16 $\mathsf{DU}$ 64 8
  17 $\mathsf{BN_{D}}$ 2 3
  18 $\mathsf{PC}$ 0 0
  19 $\mathsf{UDB}$ 258 35
  20 $\beta\mathsf{U}$ 256 16
  21 $\mathsf{DLU}$ 128 16
  22 $\mathsf{AD}$ 6 2
  23 $\mathsf{DPA\mbox{-}SNR}$ 6.312455 3.015113
  24 $\mathsf{TO}$ 7.174510 4.258065
Table 7: Results of S-box Analysis – $4\times 4$ bit case
  S# Properties
New Sbox
$4\times 4$
Elephant
$4\times 4$
GIFT-COFB
$4\times 4$ GIFT
Photon-Beetle
$4\times 4$ Present
  1 Bijectivity
  2 Balancedness
  3 Permutation
  4 $\mathsf{OP}$ 7 13 9 7
  5 $\mathsf{FP}$ 3 0 0 0
  6 $\mathsf{OFP}$ 1 1 1 1
  7 $\mathsf{BIC}$ 0.577350 1.000000 1.000000 1.000000
  8 $\mathsf{SAC}$ 0.7500 1.000000 1.000000 1.000000
  9 $\barwedge_{S}$ 8 16 16 16
  10 $\doublebarwedge_{S}$ 640 1024 1024 1024
  11 $\xi$ 0.25000 0.25000 0.25000 0.25000
  12 $\mathsf{NL}$ 4 4 4 4
  13 $\mathsf{BN_{L}}$ 2 2 2 2
  14 $\mathsf{LS}$ 0 9 9 9
  15 $\mathsf{CI}$ 0 0 0 0
  16 $\mathsf{DU}$ 4 4 6 4
  17 $\mathsf{BN_{D}}$ 2 3 2 3
  18 $\mathsf{PC}$ 0 0 0 0
  19 $\mathsf{UDB}$ 0 3 6 6
  20 $\beta\mathsf{U}$ 10 16 16 16
  21 $\mathsf{DLU}$ 4 8 8 8
  22 $\mathsf{AD}$ 3 3 3 3
  23 $\mathsf{DPA\mbox{-}SNR}$ 2.578634 2.398501 2.398501 2.128608
  24 $\mathsf{TO}$ 3.400000 3.266667 3.466667 3.533333
3 Cryptographic Properties of S-boxes and their Classification
The strength of cryptographic primitives is best analysed with the help of well-known cryptanalysis techniques. Shannon’s theory of confusion and diffusion plays the foundational role in the design and analysis of iterative ciphers. The aspect of confusion is introduced in symmetric ciphers using highly non-linear components such as S-boxes. Designers and cryptanalysts treat these ciphers as a white box, analysing all the underlying mathematical components in order to understand and rule out any potential attack vectors, by imposing design practices based on standard characteristics and theoretically well-defined bounds. Owing to the importance of S-boxes in the design of cryptographic algorithms, an extensive research has been done over the years to explain the most refined design and analysis criteria that could be employed in order to secure them against potential attack vectors. These properties can be classified in a number of ways depending whether upon the types of representation in boolean algebra or the statistical and implementation based attacks that could be launched through exploitation of relevant metrics or via studying the inter-dependencies that link one property to another.
This section briefly covers an introduction to around twenty-four cryptographic properties of S-boxes that hold significance in the current literature. Each of the properties is related to one or more crytanalysis techniques, that could exploit the non-linear part of a block cipher, see Table 3. In addition, Table 4 presents the upper and lower bounds (UB & LB) on these properties with respect to the dimension of S-box being used [88, 89, 71, 90, 91, 92, 72, 73]. We also talk about the behaviour of these properties under affine equivalence, see Table 5. The properties along with their respective bounds and equivalence characteristics aid designers and cryptanalysts in the selection of good S-boxes.
3.1 Diffusion Characteristics
Properties such as balancedness, bijectivity, order of permutation, fixed and opposite fixed points, strict avalanche and bit independence criteria, absolute and sum of square indicator are general pre-requisites that should adhere to defined bounds for any S-box in order to avoid correlations between the input and output bits and serve as a pre-requisite to ensure security against various statistical cryptanalysis attacks.
Balancedness
A vectorial boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ is balanced if it has an equal number of $1$’s and 0’s in its coordinate and component functions. Mathematically;
$P(0)=P(1)=\frac{1}{2}=50\%$ (1)
An imbalanced boolean mapping stays at a risk of achieving high linear approximation probabilities [93]. Balancedness is invariant when under affine equivalence, which means that the number of 0’s ad $1$’s is preserved in $S$ if an affine transformation is applied to it [94, 95, 96].
Permutation
Order of Permutation (OP)
Fixed Points (FP)
$S(\varkappa)=\varkappa$ (2)
Opposite Fixed Points (OFP)
$S(\varkappa)=({\varkappa})^{\prime}$ (3)
Bit Independence Criteria (BIC)
[…]
Where; $u\neq v\$and u,v=1,2,..,m\ u,v=1,2,.....,m\ . The numerator in eq 4 denotes co-variance between two avalanche variables ($\gamma_{u}$ and $\gamma_{v}$), while $\sigma(\gamma_{u})$ and $\sigma(\gamma_{u})$ in the denominator is the product of the same variables. For the coordinate and component functions of $S$, the BIC value should stay in the range of $[0,1]$. An ideal S-box has a BIC value equal to 0. If an affine transformation is applied to $S$, BIC does not stay invariant [100].
@@ 3.1 Diffusion Characteristics
Where $\gamma\in F_{2}^{n},\gamma\neq 0,wt(\gamma)=1$. SAC is variant under affine transformation, which means that a function not satisfying SAC could be transformed to a function that satisfies SAC using affine equivalence [101].
@@
The overall impact of auto correlation of a boolean function or a vectorial boolean function is studied by finding the maximum value in ACT spectrum, known as absolute indicator (AI). Let $f$ be a Boolean function over $F_{2}^{n}$ i.e. $f:F_{2}^{n}\rightarrow F_{2}$. The AI of $f$ is defined as [88]:-
@@
Here $D_{\gamma}f=f(\varkappa)\oplus f(\varkappa\oplus a)$.
@@
$\barwedge_{S}$ stays invariant under affine transformation. For balanced vectorial boolean functions, the value of AI is always 0. If a boolean mapping has high $\barwedge_{S}$, it becomes highly susceptible to cube attacks [103, 88].
@@
SSI is not variant under affine equivalence [104]. For achieving good diffusion characteristics in a cryptographic algorithm, S-boxes should have low SSI and AI values [105, 88].
@@ 3.2 Linear Cryptanalysis
The idea of this technique evolves around studying the non-linear layer of a block cipher to draw a linear approximations profile (LAT) between plaintext, key and the ciphertext to find correlations among them. Matsui introduced non-linearity as a significant criteria in the design of block ciphers [106]. Block ciphers having a highly non-linear layer are resistant towards linear cryptanalysis. To study the number of active S-boxes in the linear trails of a block cipher, J. Daemon et al. [107] introduced linear branch number. Theoretically, in order to resist linear (and its variant) cryptanalysis techniques an S-box should have a low linear approximation probability, high non-linearity, high order of linear branch number, equal to none linear structures, and low order correlation immunity.
@@
$\mathsf{LAT}(\varrho,\psi)=|\{\#\varkappa\ |\ \varrho\cdot\varkappa=\psi\cdot S(\varkappa)\}|-2^{n-1}$ (11)
Definition 1.
@@
$\xi=\frac{1}{2^{n}}\max_{\varrho,\psi\in F_{2}^{n},F_{2}^{m}\backslash\{0\}}\bigg|\mathsf{LAT}(\varrho,\psi)\bigg|$ (12)
@@
In block cipher design strategy, branch number is a significant property used to measure the strength of an S-box against linear cryptanalysis. An S-box with high branch number could increase the diffusion power of a cryptographic algorithm. Mathematically:-
@@
LBN is variant under affine transformation i.e. two S-boxes can be affine equivalent of each other but they can have different values of LBN [112].
@@
S-boxes having LS are vulnerable to attacks more robust than exhaustive key search, such as differential and linear cryptanalysis [114]. LS is an affine variant property of any boolean mapping $S$ [115, 88].
@@
$\{\mathsf{W}(\gamma,\rho)=0:\ \gamma\in F_{2}^{m}\backslash\{0\},\ \rho\in F_{2}^{n},\ \forall\ 1\leq wt(\rho)\leq k\}$ (17)
CI shows variant behavior when an affine transformation is applied to $S$ [88]. For a good cryptographic S-box the lower is the order of CI the better. CI shares an inverse relation with NL, hence if CI holds a lower order the value of NL will be high, and vice versa.